OpenStudy (anonymous):

integral from 0 to infinity of: lnx/x^2

OpenStudy (anonymous):

do you mean \[ \frac{ \ln x }{ x ^{2} }\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i would say parts for this one to find the anti derivative i mean

OpenStudy (misty1212):

tru \(u=\ln(x), dv =\frac{1}{x^2}\)

OpenStudy (anonymous):

parts looks good

OpenStudy (anonymous):

cause the next one will be a u-sub

OpenStudy (idku):

\[\int\limits_{ }^{ } \frac{\ln(x)}{x^2}~dx\]\[u=\ln(x)~~~~~~~~~du=1/x~dx~~~~~~~~~~~1/x=1/e^u\]\[\int\limits_{ }^{ } \frac{u}{e^u}~du\]and then do parts.

OpenStudy (idku):

that is my first look at it.

OpenStudy (idku):

\[v=\int\limits_{ }^{ } e^{-u}~du=-e^{-u}\]\[u=u,~~~~>>~~du=1\]

OpenStudy (idku):

\[\int\limits_{ }^{ } \frac{u}{e^u}~du=-\frac{u}{e^u}+\int\limits_{ }^{ }e^{-u}~du\]

OpenStudy (idku):

then it should not be hard... after you have the answer in terms of u, sub the x back in

OpenStudy (thomas5267):

Why not by parts though? By parts is easy...

OpenStudy (anonymous):

where did the e come from?

OpenStudy (idku):

I am using by parts, but I make mit a bit simpler by u sub on the first step, no?

OpenStudy (idku):

sateline, what do you mean?

OpenStudy (thomas5267):

No you are making it more complex I think.

OpenStudy (idku):

it is easier to integrate e^(-u) and to differentiate the u. and my sub was not that difficult.

OpenStudy (anonymous):

you get \[-\frac{\ln(x)}{x}+\int \frac{dx}{x^2}\] in one step

OpenStudy (idku):

well... actually, yes. I change my mind.

OpenStudy (idku):

it is easier to do by parts from the begin.

OpenStudy (idku):

ln(x)/x is the thing that made me think of a u sub.... it is longer, but not a bad exercise.